# LoRA Learns Less and Forgets Less

### v1

@@ 4.2 LoRA forgets less than full finetuning
Figure 3: LoRA forgets less than full finetuning. We plot the average of HellaSwag, ARC-Challenge and Winogrande for Llama-2-7B trained trained on: (A) StarCoder-Python (B) Magicoder-Evol-Instruct-110k (C) OpenWebMath (D) MetaMathQA.
We define forgetting as the degradation in the average of HellaSwag, ARC-challenge, and WinoGrande benchmarks, and investigate its extent as a function of data in Fig. 3.
Overall, we observe that (1) IFT induces more forgetting than than CPT, (2) programming induces more forgetting than math, and (3) forgetting tends to increase with data. Most importantly, LoRA forgets less than full finetuning, and as in 4.1, the effects are more pronounced for the programming domain. In code CPT, LoRA’s forgetting curve is roughly constant, whereas full finetuning degrades with more data (the forgetting metric at peak HumanEval: Full finetuning=0.54 at 20B tokens, LoRA=0.64 at 16B tokens). In programming IFT, both methods degrade when trained for more epochs, and at their peak performance (4 and 8 epochs), LoRA scores 0.63 and full finetuning scores 0.45. For math, there are no clear trends on the OpenWebMath CPT dataset, except that both LoRA and full finetuning exhibit no forgetting. This is likely due to the fact that the OpenWebMath dataset is dominated by English sentences, unlike the StarCoder-Python dataset which is majority Python code (see 3.1 for details). In math IFT, LoRA again forgets less than full finetuning (0.63 versus 0.57, repectively, at epoch 4).
Figure 4: LoRA vs. Full Finetuning trade-off for LLaMA-2-7B. Relative to full finetuning, LoRA learns less (lower values on the $y$-axis) and forgets less (higher values on the $x$-axis). Dots represent individual models trained for various epochs. For LoRA models, each configuration is shown as a separate dot. In panel B, we scatter four additional full finetuning models with non-zero attention dropout and weight decay, showing epochs 1,2,4 and 8. Same data as Figures 2, 3 and S5.
@@ 4.3 The Learning-Forgetting Tradeoff
It is trivial that models that change less when finetuned to a new target domain, will forget less of the source domain. The nontrivial question is: do LoRA and full finetuning differ in how they tradeoff learning and forgetting? Can LoRA achieve similar target domain performance but with diminished forgetting?
We form a learning-forgetting Pareto curve by plotting the aggregate forgetting metric vs. the learning metric (GSM8K and HumanEval), with different models (trained for different durations) scattered as points in this space (Fig. 4). LoRA and full finetuning seem to occupy the same Pareto curve, with the LoRA models on the lower right – learning less and forgetting less. However, we are able find cases, especially for code IFT, where for a comparable level of target-domain performance, LoRA exhibits higher source-domain performance, presenting a better tradeoff. In supplementary Fig. S5 we show the raw evaluation scores for each model. In Fig. S3 we scatter the Llama-2-13B results in the same plot as Llama-2-7B for Code CPT.
Figure 5: LoRA provides stronger regularization compared to attention dropout and weight decay. LoRA finetuning (green) leads to less learning (as measured by accuracy on HumanEval, left) and less forgetting (as measured by HellaSwag, ARC and WinoGrande, right).
4.4 LoRA’s regularization properties
Here, we define regularization (loosely) as a training mechanism that keeps the finetuned LLM similar to the base LLM. We first analyze similarity in the learning-forgetting tradeoff and then in the generated text.
@@ 4.4 LoRA’s regularization properties
LoRA is a stronger regularizer compared to weight decay and attention dropout.
We use Llama-2-7B models trained on Magicoder-Evol-Instruct-110K dataset, and compare LoRA ($r=16,256$, “All”) to weight decay (Goodfellow et al., 2016) with values $5e^{-5},1e^{-4}$ and attention dropout (Srivastava et al., 2014) with values $0.05,0.1$. We find that LoRA provides stronger regularization: it learns less and forgets less.
@@
We again use Llama-2-7B models trained on Magicoder-Evol-Instruct-110K dataset to scrutinize the generated strings during HumanEval. We calculate the unique number of output strings out of 50 generations (for base model, full finetuning, and LoRA) serving as a coarse proxy for predictive diversity. In Figure 6 we separately show the results for correct and incorrect answers. As in the reinforcement learning from human feedback literature (Du et al., 2024; Sun et al., 2023), we find that full finetuning results in fewer unique generations (“distribution collapse”) compared to the base model, for both pass and fail generations. We find that LoRA provides a compromise between the two, at the level of generations. The above works also suggest that LoRA could even substitute a common regularization term that keeps the probabilities of the generated text similar between the finetuned and base model.
@@ 4.5 Full finetuning on code and math does not learn low-rank perturbations
First, in Figure S6 we present results for the $W_{q}$ projection at layer 26 of Llama-2-7B (with dimensions $4096\times 4096$). We show that the spectrum of the finetuned weight matrix is very similar to that of the base weight matrix, both decaying slowly and requiring keeping $\approx 50\%$ of singular vectors ($\approx 2000/4096$) to explain 90% of the variance in the weight matrix. Critically, the difference $\Delta$ also has a similar spectrum to the finetuned and base weight matrices (up to a multiplicative scaling). We suggest that there is nothing extraordinary about the full finetuning spectra; similar spectra can be achieved by adding low-magnitude Gaussian i.i.d noise to a weight matrix (Fig. S7).
@@
Next, we ask when during training does the perturbation become high rank, and whether it meaningfully varies between module types and layers. We estimate the rank needed to explain 90% of the variance in the matrix. The results appear in Figure 7. We find that: (1) The earliest checkpoint at 0.25B CPT tokens exhibits $\Delta$ matrices with a rank that is 10-100X larger than typical LoRA ranks; (2) the rank of $\Delta$ increases when trained on more data; (3) MLP modules have higher ranks compared to attention modules; (4) first and last layers seem to be lower rank compared to middle layers.
@@
Figure 8: LoRA is more sensitive to learning rates compared to full finetuning. LLaMA-2-7B models (A) trained on Magicoder-Evol-Instruct-110k (Wei et al., 2023) and evaluated on HumanEval, (B) trained on MetaMathQA (Yu et al., 2023) and evaluated on GSM8K. Experiments here are performed with LionW; see Fig. S1 for a comparion to AdamW.
(a) LLaMA-2-7B trained on Magicoder-Evol-Instruct-110K for up to 16 epochs.
(b) LLaMA-2-7B trained on MetaMathQA for up to 16 epochs.
Figure 9: The improvement from targeting more modules (shown as different colors) is more substantial than from increasing the rank by 16$\times$ ($x$-axis). Different points indicate different training durations.
4.6 Practical takeaways for optimally configuring LoRA
Though optimizing LoRA hyperparameters does not close the gaps with full finetuning, some hyperparamater choices are substantially more effective than others, as we highlight below.
4.6.1 LoRA is highly sensitive to learning rates
We analyze learning rate sensitivity for Llama-2-7B, trained for two epochs on the code and math IFT datasets, and followed by HumanEval and GSM8K evaluation, respectively. For LoRA, we set $r=64$, and target modules$=$“All”. Fig. 8 shows that LoRA improves monotonically with learning rate up to a value at which training diverges, with best learning rates of $5e^{-4}$ for code and $2e^{-4}$ for math. On both datasets, these best LoRA learning rates are underperformed by four alternative full finetuning learning rates. The best full finetuning learning rates are $5e^{-5}$ and $1e^{-5}$, respectively, an order of magnitude smaller than LoRA. For LoRA, we cannot find alternative learning rates that achieve at least 90% of the best learning rate’s performance. For full finetuning, there are two viable alternative learning rates for code and three for math.
4.6.2 Choice of target modules matters more than rank
With the best learning rates at hand, in Fig. 9, we proceed to analyze the effect of rank ($r=16,256$) and target modules. We find that “All” $>$ “MLP” $>$ “Attention” and that though the effects of rank are more subtle, $r=256>r=16$. We therefore conclude that targeting “All” modules with a relatively low rank (e.g., $r=16$) provides a good tradeoff between performance and accuracy.
All in all, we recommend using LoRA for IFT and not CPT; identifying the highest learning rate that enables stable training; targeting “All” modules and choosing rank according to memory constraints, with 16 being a good choice; exploring training for at least four epochs.
@@ 5 Related Work
LoRA has inspired many variants and extensions that are more memory-efficient or performant, such as QLoRA (Dettmers et al., 2024), VeRA (Kopiczko et al., 2023), DoRA (Liu et al., 2024), Chain of LoRA (Xia et al., 2024), and GaLoRe, (Zhao et al., 2024a), as well as efficient inference techniques building on LoRA, such as S-LoRA (Sheng et al., 2023).
@@
The original LoRA study by Hu et al. (2021) reported that LoRA matched full finetuning performance for RoBERTa (Liu et al., 2019) on GLUE (Wang et al., 2018), and GPT-2 on E2E NLG Challenge (Novikova et al., 2017), and GPT-3 on WikiSQL (Zhong et al., 2017), MNLI (Williams et al., 2017), and SAMSum (Gliwa et al., 2019). Many subsequent studies follow this template and report encoder model performance on tasks in GLUE such as SST-2 (Socher et al., 2013) and MNLI (Williams et al., 2017). Models such as RoBERTa are less than 340M parameters, however, and classification tasks such as MNLI are quite trivial for modern billion-parameter LLMs such as Llama-2-7B. Despite LoRA’s popularity, only a few studies have rigorously compared LoRA to full finetuning in this setting and with challenging domains such as code and math. Dettmers et al. (2024) for example found that QLoRA matched full finetuning MMLU (Hendrycks et al., 2020) performance when finetuning Llama-1 7B, 13B, 33B and 65B on the Alpaca (Taori et al., 2023) and FLAN (Chung et al., 2024) datasets. Ivison et al. (2023) on the other hand found that QLoRA did not perform as well as full finetuning for Llama-2-7B, 13B and 70B models trained on the Tülü-2 dataset when evaluated across a suite of tasks including MMLU, GSM8K, AlpacaEval (which uses LLM-as-a-judge; (Dubois et al., 2024)) and HumanEval. One recent notable study is Astraios, which found that LoRA performed worse than full finetuning on 8 datasets and across 4 model sizes (up to 16 billion parameters), on 5 representative code tasks (Zhuo et al., 2024). Our study corroborates these results.
@@
The conclusions have also been mixed with regards to the practical details surrounding LoRA target modules and rank: Raschka (2023) and Dettmers et al. (2024) show that optimized LoRA configurations perform as well as full finetuning, and that performance is governed by choice of target modules but not rank. In contrast, Liu et al. (2024) shows that LoRA is sensitive to ranks. It is likely that some of these discrepancies are due to differences in finetuning datasets and evaluations.
@@
Continual learning on a new target domain often comes at the expense of performance in the source domain (Lesort et al., 2020; Wang et al., 2024). A relevant example is that code-finetuned LLMs lose some of their capabilities in language understanding and commonsense reasoning (Li et al., 2023; Roziere et al., 2023; Wei et al., 2023). A common approach to mitigate forgetting involves “replaying” source-domain data during continual learning, which can be done by storing the data in a memory buffer, or generating it on the fly (Lesort et al., 2022; Scialom et al., 2022; Sun et al., 2019).

### v2

@@ 4.2 LoRA forgets less than full finetuning
Figure 2: LoRA forgets less than full finetuning. In all panels, the $y$-axis shows the average of HellaSwag, ARC-Challenge and Winogrande for Llama-2-7B trained trained on: (A) StarCoder-Python (B) Magicoder-Evol-Instruct-110k (C) OpenWebMath (D) MetaMathQA.
Here, we investigate the extent of forgetting (defined in Sec. 3.2) as a function of training data in Fig. 2.
Overall, we observe that (1) IFT induces more forgetting than than CPT, (2) programming induces more forgetting than math, and (3) forgetting tends to worsen with training duration. Most importantly, LoRA forgets less than full finetuning, and the extent of forgetting is controlled by rank. In code – for both CPT and IFT – full finetuning forgets substantially more than any LoRA configuration. In code CPT (Table S2), at 20B tokens, full finetuning scores 0.545 versus 0.617 by LoRA $r=256$. In code IFT (Table S6), full finetuning scores 0.414 versus 0.509 by LoRA $r=64$. In math – for both CPT and IFT – LoRA with $r=256$ forgets nearly as much as full finetuning. In CPT (Table S4), LoRA scores 0.616 (20B tokens) versus 0.613 of full finetuning (16B tokens). In IFT (Table S8), LoRA and full finetuing respectively degrade to 0.567 and 0.559 at epoch 16.
We note that the least forgetting occurs for the OpenWebMath dataset, which is dominated by English sentences (see 3.1 for details).
Figure 3: LoRA vs. full finetuning tradeoff for Llama-2-7B. Relative to full finetuning, LoRA learns less (lower values on the $y$-axis) and forgets less (higher values on the $x$-axis). Each dot is a separate model, with marker size corresponding to training duration (from 0.25-20 billion tokens for CPT, and 1-16 epochs for IFT). Same data as Figures 1, 2.
@@ 4.3 The Learning-Forgetting Tradeoff
It is trivial that models that change less when finetuned to a new target domain will forget less of the source domain. The nontrivial question is: do LoRA and full finetuning differ in how they trade off learning and forgetting? Can LoRA achieve similar target domain performance but with diminished forgetting?
We form learning-forgetting Pareto curves by plotting the forgetting metric versus the learning metric for each training duration (Fig. 3). As models train on more data, they learn more and forget more, traveling up and left in this space. As we increase LoRA ranks, we find that the curves shift up and left as well, again, learning more and forgetting more, doing so more consistently in IFT than CPT.
Each dataset presents a unique tradeoff pattern which makes it difficult to conclude whether LoRA and full finetuning offer fundamentally different learning-forgetting tradeoffs. We will review each dataset next.
For Code CPT, though the full finetuning curve reaches much higher values of HumanEval, it appears to forget more for any given HumanEval value, which LoRA can reach if trained on more tokens. This pattern does not hold for math CPT, where LoRA and full finetuning curves are roughly overlapping until full finetuning shoots up (in 4B tokens) to achieve much higher GSM8K scores without increased forgetting. In code IFT, LoRA $r=256$ offers comparable HumanEval accuracy while strictly forgetting less. Lower ranks do not reach high values on HumanEval to compare to full finetuning. In math IFT, LoRA and full finetuning seem to lie on adjacent learning-forgetting tradeoff curves, with full finetuning offering preferable tradeoffs.
With the caveats mentioned above, it seems that LoRA can offer preferable learning-forgetting tradeoffs for code, while full finetuning can offer preferable tradeoffs for math. Moreover the choice of LoRA rank can serve as a knob to navigate the learning-forgetting tradeoffs.
Figure 4: LoRA forgets less than attention dropout and weight decay. Results from Llama-2-7B finetuned on Magicoder-Evol-Instruct-110K. Left panel: learning as measured by accuracy on HumanEval. Right panel: forgetting as measured by the average of HellaSwag, ARC-Challenge and WinoGrande scores. The solid slateblue line shows that LoRA (r=256) learns as much as full finetuning, weight decay, and attention dropout, while forgetting much less.
4.4 For the Tülu-v2-mix dataset, LoRA is on par with full finetuning
So far, we analyzed how LoRA and full finetuning specialize in very specific domains. Often, code or math problems appear as part of larger IFT data mixtures that include multi-turn conversations and a variety of other NLP tasks, such as summarization, etc. (e.g. Wei et al. (2021)). We therefore finetuned LoRA and full finetuning models on one such popular dataset, the Tülu-v2-mix (Ivison et al., 2023). The results are presented in the Appendix (Sec. C and Table S13). In summary, we find that both LoRA and full finetuning meaningfully improve upon the base model, and that LoRA, even with lower ranks, can match full finetuning in chat quality as measured by Multi-Turn Benchmark (MT-bench (Zheng et al., 2024)), GSM8K (Cobbe et al., 2021), and Massive Multitask Language Understanding (MMLU; Hendrycks et al. (2020)). At longer training durations (6 epochs), LoRA also forgets less.
4.5 How strongly does LoRA constrain the finetuning process?
In this section, we analyze Llama-2-7B models trained on the Magicoder-Evol-Instruct-110K dataset. We first compare the learning-forgetting tradeoffs between LoRA and classic regularization techniques, and then analyze the diversity of the generated text.
@@ 4.5 How strongly does LoRA constrain the finetuning process?
LoRA forgets less than attention dropout and weight decay
We compare LoRA ($r=16,256$, training all modules) to weight decay (Goodfellow et al., 2016) with values $5e^{-5},1e^{-4}$ and attention dropout (Srivastava et al., 2014) with values $0.05,0.1$. Both regularization techniques appear to learn and forget as much as full finetuning, except that weight decay starts to generally deteriorate at longer training durations (epochs 8 and 16). LoRA, with the common $r=16$, learns less and forgets less than all other models. LoRA $r=256$, on the other hand, learns as much as the other methods while forgetting less.
@@
We scrutinize the generated solution strings for HumanEval problems. We calculate the unique number of output strings out of 50 generations (for base model, full finetuning, and LoRA) serving as a coarse proxy for predictive diversity. In Figure 5 we separately show the results for correct and incorrect answers. As in the reinforcement learning from human feedback literature (Du et al., 2024; Sun et al., 2023), we find that full finetuning results in fewer unique generations (“distribution collapse”) compared to the base model, for both pass and fail generations, with LoRA in between the two. The above works also suggest that LoRA could even substitute a common Kullback-Leibler divergence term that keeps the probabilities of the generated text similar between the finetuned and base model. We reiterate that exact string matching between generations is not a sensitive metric of predictive diversity, as generations can slightly vary in format and remain functionally identical.
@@ 4.6 Full finetuning on code and math does not learn low-rank perturbations
First, in Figure S7 we present results for the $W_{q}$ projection at layer 26 of Llama-2-7B (with dimensions $d\times d$, $d=4096$). We show that the spectrum of the finetuned weight matrix is very similar to that of the base weight matrix, both decaying slowly and requiring keeping $\approx 50\%$ of singular vectors ($\approx 2000/4096$) to explain 90% of the variance in the weight matrix. Critically, the difference $\Delta$ also has a similar spectrum to the finetuned and base weight matrices (up to a multiplicative scaling). These results are in line with the analysis in Zeng & Lee (2024) showing that any transformer model can be well approximated with $r=d/2$. Additionally, we suggest that there is nothing extraordinary about the full finetuning spectra; similar spectra can be achieved by adding low-magnitude Gaussian i.i.d noise to a weight matrix (Fig. S8).
@@
Next, we ask when during training does the perturbation become high rank, and whether it meaningfully varies between module types and layers. We estimate the rank needed to explain 90% of the variance in the matrix. The results appear in Figure 6. We find that: (1) The earliest checkpoint at 0.25B CPT tokens exhibits $\Delta$ matrices with a rank that is 10100×10-100$\times$10 - 100 × larger than typical LoRA ranks; (2) the rank of $\Delta$ increases when trained on more data; (3) MLP modules have higher ranks compared to attention modules; (4) first and last layers seem to be lower rank compared to middle layers.
@@
Figure 7: Targeting MLP or All modules is superior to training Attention modules alone. All Llama-2-7B checkpoints were trained on Magicoder for 1, 2 and 4 epochs with rank 16 (left), 64 (center) and 256 (right).
4.7 Hyperparameter sensitivity analyses for LoRA
Our goal in this work was to optimally configure LoRA so that it has the best chances of matching full finetuning. This is nontrivial, as LoRA has a large number of hyperparameters to choose from: target modules, rank, scaling factors, and learning rates. We turn to analyze the importance of each, and provide some practical recommendations.
First, we found that the choice $\alpha=2r$ is crucial for high ranks. Most common packages, e.g. HuggingFace’s peft, [^11] https://huggingface.co/docs/peft/en/index scale the LoRA matrices by $\alpha/r$, effectively scaling down higher ranks (see also Kalajdzievski (2023)). One might think that high learning rate values may compensate for fixed low $\alpha$’s, but doing so creates instabilities and often leads to inferior performance. To show this, we performed a joint hyperparameter sweep over $\alpha$ and learning rate for the Magicoder dataset training a $r=256$ LoRA for 4 epochs (Fig. S3). We find that $\alpha=512$ does much better than $256$ or $32$ across all learning rates.
Next, to assess the relative contribution of target modules and rank, we trained Llama-2-7B models on 4 epochs of the Magicoder dataset, sweeping over target modules (“Attention”, “MLP”, and “All”, their union), ranks ($r=16,64,256$), setting $\alpha=2r$. Fig. 7 shows that HumanEval performance increases with rank, and that targeting just “Attention” underperforms both “MLP” and “All”, where in the latter, most gains are interestingly driven by the “MLP” modules. This is potential evidence that the MLP blocks are the primary loci for continual learning in LoRA, at least in our datasets.
For IFT, we find that LoRA is more sensitive to learning rates compared to full finetuning, and benefits from the highest learning rate that enables stable training for the chosen training duration (see Appendix Sec. B and Fig. S1). LoRA’s best learning rates should be set one order of magnitude higher than that of full finetuning, often ranging between $5e^{-5}$ and $5e^{-4}$ for these combinations of model architecture and dataset.
In Appendix Sec. I, we benchmark throughput and peak GPU memory of different LoRA configurations, showing that for standard implementations and a fixed batch size, LoRA tends to train slower than full finetuning.
To conclude, based on our main results and hyperparameter sweeps, we recommend: (a) using LoRA for instruction finetuning and not continued pretraining; (b) if GPU memory allows, targeting “All” transformer modules with a rank of $256$, since ranks $16-64$ tend not to suffice for code tasks; (c) using $\alpha=2r$, and (d) sweeping over learning rates between $[1e-5,5e-4]$, picking the highest value that enables stable training.
@@ 5 Related Work
LoRA has inspired many variants and extensions. One group of methods improves training with LoRA by focusing on initialization or scaling (Meng et al., 2024; Hayou et al., 2024; Li et al., 2023b; Kalajdzievski, 2023; Nikdan et al., 2024), sequential training procedures (Xia et al., 2024), or architectural modifications (Shi et al., 2024). Other works propose alternative low-rank approximations altogether (Liu et al., 2024; Zhao et al., 2024a; Jiang et al., 2024a; Kopiczko et al., 2023). In this study we chose to analyze the classic LoRA setup; while many of these proposed variations of LoRA seem promising, we leave a rigorous comparison of these techniques to future work.
@@
The original LoRA paper Hu et al. (2021) reported that LoRA matched full finetuning performance for RoBERTa (Liu et al., 2019) on GLUE (Wang et al., 2018), GPT-2 on E2E NLG Challenge (Novikova et al., 2017), and GPT-3 on WikiSQL (Zhong et al., 2017), MNLI (Williams et al., 2017), and SAMSum (Gliwa et al., 2019). Many subsequent studies follow this template and report encoder model performance on tasks in GLUE such as SST-2 (Socher et al., 2013) and MNLI (Williams et al., 2017). Models such as RoBERTa are less than 340M parameters, however, and classification tasks such as MNLI are quite trivial for modern billion-parameter LLMs such as Llama-2-7B. Despite LoRA’s popularity, only a few studies have rigorously compared LoRA to full finetuning in this setting and with challenging domains such as code and math. Dettmers et al. (2024) for example found that QLoRA matched full finetuning MMLU (Hendrycks et al., 2020) performance when finetuning Llama-1 -7B, 13B, 33B and 65B on the Alpaca (Taori et al., 2023) and FLAN (Chung et al., 2024) datasets. Ivison et al. (2023) on the other hand found that QLoRA did not perform as well as full finetuning for Llama-2-7B, 13B and 70B models trained on the Tülü-v2-mix dataset when evaluated across MMLU, GSM8K, AlpacaEval (which uses LLM-as-a-judge; (Dubois et al., 2024)) and HumanEval. One recent notable study is Astraios, which found that LoRA at rank $r=8$ performed worse than full finetuning on 8 datasets and across 4 model sizes (up to 16 billion parameters), on 5 representative code tasks (Zhuo et al., 2024). Our study corroborates these results and shows that with higher ranks and proper hyperparameter choices, LoRA can perform much better.
@@
The conclusions have also been mixed with regards to the practical details surrounding LoRA target modules and rank: Raschka (2023) and Dettmers et al. (2024) show that optimized LoRA configurations perform as well as full finetuning, and that performance is governed by choice of target modules but not rank.121212see also Zhang et al. (2024), who report some cases where performance does improve with rank. However, in that work, the scalar $\alpha$ was not modified with rank, and we found that increasing it to $2r$ was necessary to unlock improvements by rank. In contrast, Liu et al. (2024) shows that LoRA is sensitive to ranks. It is likely that some of these discrepancies are due to differences in finetuning datasets and evaluations.
Continual learning on code and math.
A growing body of work investigates ways of specializing LLMs for code and math. In code, models such as StarCoder (Li et al., 2023a; Lozhkov et al., 2024), DeepSeek Coder (Guo et al., 2024), and SantaCoder (Allal et al., 2023) were pretrained from scratch on large-scale code datasets. Alternatively, some works start with a generic pretrained base model, and combine continued pretraining on large code datasets followed by IFT on code problems (usually with full finetuning), e.g., Codex (Chen et al., 2021), Code-Qwen (Bai et al., 2023), CodeLlama (Roziere et al., 2023). Some perform only IFT on top of a base model, like MagiCoder (Wei et al., 2023), or WizardCoder (Luo et al., 2023b). Other models such as OctoCoder (Muennighoff et al., 2023) perform IFT with LoRA.
Similarly, much recent work aims to improve mathematical capabilities. Models like DeepSeek Math (Shao et al., 2024) perform continued pretraining on top of a base model, while other methods focus on finetuning by generating high-quality synthetic math problems, scaling to millions of examples. Luo et al. (2023a) takes the Evol-Instruct approach to data generation (akin to the Magicoder dataset; Sec. 3.1) which it then uses to train reward models for instruction quality and solution correctness, which are in turn used for LLM finetuning. Other work develops Monte Carlo Tree Search methods to automatically supervise the intermediate reasoning steps while solving math problems (Luo et al., 2024), and Yue et al. (2024) generates questions and answers from the pretraining web corpus. Toshniwal et al. (2024) uses an LLM to synthesize Code-Interpreter-style solutions to the GSM8K and MATH benchmarks; the proposed solutions can be verified against the official solutions. Singh et al. (2023) iterate over this procedure multiple times (“Self-training”) using an expectation-maximization approach. All reviewed methods meaningfully improve math capabilities.
@@
Vu et al. (2022) shows that prompt tuning (Lester et al., 2021), another parameter-efficient finetuning method, can aid in mitigating forgetting for cross-lingual summarization tasks (using multilingual variants of the T5 model). With large Llama-style LLMs, it has been reported that code-finetuned LLMs lose some of their capabilities in language understanding and commonsense reasoning (Li et al., 2023a; Roziere et al., 2023; Wei et al., 2023). A common approach to mitigate forgetting involves “replaying” source-domain data during continual learning, which can be done by storing the data in a memory buffer, or generating it on the fly (Lesort et al., 2022; Scialom et al., 2022; Sun et al., 2019).